# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ III.1 Nuclear Structure Functions and nPDFs
In 1981, the EMC collaboration measured the nuclear modification of structure functions in deep inelastic scattering (DIS) [1]. Due to the relation between structure functions and parton distribution functions at leading order and leading twist,
@@
these data revealed, for the first time, that the parton distribution function (PDF) of a proton bound within a nucleus (nPDF) differs from that of a free proton: $f_{i}^{p/A}(x,Q^{2})\neq f_{i}^{p}(x,Q^{2})$, where $f_{i}^{p/A}$ represents the PDF of a flavor $i$ inside a proton bound in a nucleus. The nuclear modification factor due to this effect is defined as $R_{i}^{A}\equiv f_{i}^{p/A}/f_{i}^{p}$. [^5] A slightly different definition is often adopted, $R_{i}^{A}\equiv\left(Zf_{i}^{p/A}+Nf_{i}^{n/A}\right)/\left(Zf_{i}^{p}+Nf_{i}^{n}\right)$, where $Z(N)$ is the number of protons (neutrons) in nucleus $A$, and $f_{i}^{n/A}(f_{i}^{n})$ are the bound (free) neutron PDFs. This definition takes into account the fact that the proton and neutron distributions in nuclei are always averaged and cannot be accessed separately. Since then, four main $x$ regions have been identified: at $x\gtrsim 0.7$, Fermi motion dominates with $R^{A}(x)\gtrsim 1$; in the range $0.3\lesssim x\lesssim 0.7$, the EMC region is characterized by $R^{A}(x)\lesssim 1$; between $0.1\lesssim x\lesssim 0.3$, antishadowing leads to $R^{A}(x)\gtrsim 1$; and finally, for $x\lesssim 10^{-1}$, shadowing is observed with $R^{A}(x)\lesssim 1$. Global analyses of nPDFs have been performed by the DSSZ [2], EPPS [3, 4], KSTSG [5], nCTEQ [6, 7], nNNPDF [8], and TUJU [9] collaborations. The latest analyses also include data from the LHC. For a recent review, see Ref. [208].
@@
  • Electroweak boson production: Production of $W$ and $Z$ bosons in $p+A$ collisions, $p+A\to Z,W+X$, measured at LHC at $\sqrt{s_{NN}}=5.02$ and $8.16$ TeV probe quark nPDFs and, to some extent, the gluon nPDF in the range $10^{-3}\lesssim x\lesssim 10^{-1}$ in shadowing and antishadowing regions at scale $Q^{2}\sim 10^{4}\,{\rm GeV}^{2}$ [217, 218, 219, 49, 46, 50, 52].
@@
The phenomenological extraction of the nuclear modification factor $R^{\text{nPDF}}$ critically depends on the data used in the global analyses. Earlier nPDF determinations relied primarily on DIS processes [225, 226, 227], where the initial electron interacts purely via quantum electrodynamics (QED) with the nucleus, making them ideal for probing the quark nPDFs. More recently, data in $h+A$ collisions have been incorporated to extend the range of $x$ explored and access the gluon nPDFs more directly.
Figure 5 presents a comparison of recent quark and gluon nPDFs at $Q^{2}=10$ GeV2. The EPPS21 analysis [4] incorporates the $D$ production data in their global analysis, providing a strong constraint on $R_{g}^{\text{nPDF}}$ at $x\lesssim 10^{-2}$. Their result agrees with the nCTEQ15WZ analysis [228] but deviates from NNPDF2.0 [229], which does not include these data. These observations suggest that shadowing evaluations remain highly sensitive to $h+A$ data used.
@@
Including the LHC $p+{\rm Pb}$ measurements, the variety and the kinematical reach of the data used in nPDF analyses have grown significantly. It has therefore become increasingly possible to look for deviations from collinear factorization. For example, some of the nuclear effects seen in heavy flavor production in $p+{\rm Pb}$ collisions could be due to dynamical effects such as saturation [230], partonic energy loss [231], or collective effects. It is important to consider to what extent these effects are currently being incorporated in the nPDFs and how to potentially distinguish their ”non-factorizable” share. However, if the heavy flavor data are left out of the analysis, the small-$x$ gluon nPDF would be largely unconstrained and, as a result, the uncertainty of the baseline upon which these novel effects are supposed to be observed becomes prohibitively large. In other words, the baseline uncertainties overwhelm the effect one is trying to discover. In principle, this dilemma can be resolved if the remaining data are sufficiently precise and diverse, which is currently not the case. Because the factorizable effects are universal, the presence of effects beyond the nPDF formalism must eventually manifest as the inability to reproduce all the data with a single set of nPDFs in a statistically satisfactory way. It is thus useful to also view the global nPDF analyses as a search of physics beyond collinear factorization and not something that ignores or downplays them.
@@ III.2 Saturation effects in the small-$x$ regime
At small $x$ values, the gluon density inside hadrons grows rapidly, leading to nonlinear effects that are not captured by traditional linear QCD evolution equations such as DGLAP [167, 168] or BFKL [232, 233, 234, 235]. This phenomenon, known as gluon saturation, implies that the growth of the gluon distribution eventually slows down and reaches a limiting behavior. To describe this high-density regime, CGC effective field theory has been developed. It models a high energy hadron as a coherent ensemble of soft gluons with large occupation numbers. By incorporating nonlinear gluon recombination effects, the CGC provides a more accurate and unified description of high-energy scattering processes. These nonlinear dynamics are characterized by the saturation scale $Q_{s}^{2}(x)\sim A^{1/3}\,x^{-\lambda}$, which grows exponentially as $x$ decreases. At sufficiently small $x$ (i.e., at high energies), the saturation regime becomes weakly coupled due to the semi-hard scale $Q_{s}\gg\Lambda_{\text{QCD}}$, making it amenable to perturbative QCD calculations within the CGC framework.
Figure 6 illustrates the current kinematic phase space for existing DIS and Drell-Yan data from nuclei heavier than iron, as a function of $x$ and $Q^{2}$. The saturation lines $Q^{2}(x)=Q_{s}^{2}(x)$ demarcate the smooth transition between the linear and nonlinear regimes. The optimal region for studying these non linear effects is at $x\lesssim 10^{-2}$ and $Q^{2}\sim 1\text{ GeV}^{2}$. The kinematic reach of the future EIC, which will significantly extend access to this region, is also shown.
@@ III.3 Target universality from DIS to $p+A$
On the other hand, in the kinematic regime where the center-of-mass energy is very large and $Q^{2}$ is fixed but not too large, typically of the order of the saturation scale, collinear factorization is expected to break down. In this domain, the notion of weakly-interacting partons underlying parton distribution functions becomes inadequate. Instead, one must account for the nonlinear dynamics induced by large gluon densities and resum the large logarithms of the center-of-mass energy $\sqrt{s}$—specifically, $\ln(s/Q^{2})$ in DIS and $\ln(\sqrt{s}/p_{T})$ in hadronic collisions—which are commonly written as logarithms of $1/x$. The CGC formalism, where the coupling remains weak, but high-density effects are resummed to all orders, satisfies this requirement.
@@
where $x_{T},y_{T},z_{T}$ and $r_{T}$ are transverse coordinates. In multiparticle production in DIS and in proton-nucleus collisions, all other correlators are suppressed by powers of $N_{c}^{2}$ [237]. These correlation functions satisfy the JIMWLK evolution equations [238, 239] which encode the evolution of the correlators as function of $x$ once an initial condition at $x_{0}$ is specified. The large-$Q^{2}$ and small-$x$ regimes of QCD are unified in the double limit $Q^{2}\to\infty$ and $x\to 0$, where large virtuality and parton density effects interplay.
To leading order in the large $N_{c}$ approximation, the dipole and quadrupole correlators are the only objects that appear in physical cross sections within the CGC formalism. As such, they are universal quantities that can be probed by different processes, such as DIS and forward particle production in $p+A$ collisions [240, 241]. The universality of the cross sections is because the dipole and quadrupole correlators contain the small-$x$ dynamics of the target independent of the projectile. It is then possible to extract the characteristics of the correlators from DIS and apply them to other processes, conferring the CGC formalism with predictive power. This is analogous to the collinear factorization framework, which is valid at high $Q^{2}$ but is expected to break down at very small $x$. Moreover, the universality of the Wilson line correlators mirrors that of the gluon distribution function, governed by DGLAP evolution in the collinear factorization approach.
There have been sustained efforts to investigate higher-twist corrections to collinear factorization, particularly those enhanced by the size of the nuclear system, scaling as $A^{1/3}$ [242, 150, 243, 244, 245]. [^6] It is important to recognize the role of target mass corrections in the context of higher-twist effects, important at high $x$. These corrections can also be systematically included within the collinear factorization framework [246]. More recently, analogous studies have been initiated in the high-energy (small- $x$) regime, focusing on sub-eikonal, energy-suppressed contributions which are important for both unpolarized [247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258] and polarized observables [259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269], as well as for bridging the gap between the collinear and high-energy (small- $x$) frameworks [270, 271, 272, 273, 274, 275, 276, 277].
This motivates the search for observables that can be commonly accessed in both DIS and proton–nucleus collisions, enabling a consistent characterization of the small-$x$ structure of nuclei across different probes.
@@ III.4 Observables
In this section, we explore several key observables that may distinguish between nuclear parton distribution function and gluon saturation effects. So far, the extraction of nPDFs at small $x$ has not taken other physical effects into account. Determining the relative contribution of each effect is crucial, as there is limited small-$x$ data available, $x\sim 10^{-5}$ in $h+A$ collisions, as discussed in Sec. III.1.
III.4.1 Structure functions, in particular $F_{L}$
The main feature of predictions of CGC-based observables is their characteristic $x$ dependence. While the theoretical uncertainties associated with the perturbative expansion are, in principle, controllable, additional uncertainties arise from the nonperturbative inputs required to initialize small-$x$ evolution. These inputs must be modeled. The initial condition is typically parametrized and fitted to experimental data at an intermediate $x$ value, $x_{0}\sim 10^{-2}$ [278, 279, 280, 281]. Once this initial condition is fixed, the CGC framework provides first principles predictions of the $x$ dependence of the structure functions $F_{2}$ and $F_{L}$.
The EIC will have a large kinematic window in $x$, about three orders of magnitude where $x$ is small, which should allow distinction between shadowing generated by the perturbative small-$x$ evolution of the target via CGC evolution equations and that of DGLAP evolution of the $x$-dependent parameterization of shadowing models which can generate no additional shadowing since the $x$ dependence of the two are quite different. Unfortunately, the $F_{2}$ structure function necessarily receives sizable contributions from nonperturbatively large dipoles [282] at small $x$, even at intermediate $Q^{2}$, making it difficult to distinguish perturbative contributions from nuclear shadowing. While the longitudinal structure function $F_{L}$ is under much better theoretical control, it is, however, much more difficult to measure experimentally since it requires measurements at different energies.
@@ III.4.2 Single and two-particle production in $e+A$ and $p+A$
Alongside the fully inclusive processes discussed in the previous section, saturation and shadowing also leave their imprints on more differential observables, such as single and double-inclusive particle production in DIS or forward proton-nucleus collisions.
@@
A promising channel for observing saturation is the disappearance of the away-side peak in two-particle azimuthal correlations, in $e+A$ or $p+A$ collisions at forward rapidity [284, 285, 286]. In the strict back-to-back limit, CGC calculations, expanded at leading power in the ratio between the transverse momentum imbalance and the relative dijet momentum, smoothly match to TMD factorization [287]. In particular, back-to-back dihadron production at LO probes the Weizsäcker-Williams gluon distribution while direct photon-jet correlations in forward $p+A$ collisions are sensitive to the dipole gluon TMD [287]. Beyond LO in the CGC, two-particle correlations also probe the sea quark TMD distribution [288]. This correspondence between CGC predictions and TMD factorization provides a powerful framework for addressing the question posed at the beginning of Sec. III, whether the suppression of the away-side peak is purely nonperturbative in origin (as encoded in the nonpertubative part at small transverse momenta of the TMDs or in the nPDF initial condition of Collins-Soper-Sterman evolution [289, 290, 291, 292]) or whether gluon saturation itself demonstrates the perturbative mechanism at work.
@@
There is, however, a complication to the previous general picture: when the two produced particles are back-to-back in the transverse plane, higher order QCD corrections, enhanced by large logarithms in the ratio of the relative transverse momentum to the transverse momentum imbalance, further decorrelate the away side peak by a vacuum-like Sudakov suppression factor (or Collins-Soper-Sterman evolution [289, 290, 291, 292] in the TMD language).
@@
While this complication makes interpretation of the cause of suppression more ambiguous, it is systematically included both in CGC calculations and in standard TMD factorization with increasing accuracy, enabling the search for saturation effects “on top of” the modifications caused by soft gluon radiation, present even in the absence of multiple soft scattering or nonlinear effects. The interplay between Sudakov resummation and gluon saturation/small-$x$ evolution has been considered in Refs. [293, 204, 294, 295].
@@
Thus, in discussions of two-particle azimuthal correlations, it is useful to distinguish between broadening and suppression: saturation is expected to lead to both broadening (due to multiple scattering) and suppression (due to small-$x$ radiation and nonlinear evolution) whereas radiative corrections associated with large-angle soft gluon emissions (the Sudakov effect) primarily result in broadening. The CGC approach provides a physical mechanism, gluon saturation, to generate the dependence of the nuclear suppression on the $p_{T}$ of the associated hadron. In the nPDF paradigm, in turn, the $p_{T}$ dependence enters through the average $x$ value probed, the smaller the $p_{T}$, the smaller the $x$, as well as through the value of the factorization scale, typically proportional the the hadron $p_{T}$. In addition, DGLAP evolution tends to reduce the nuclear effects at higher factorization scales.
@@
Numerical predictions based on the TMD framework supplemented by CGC models if the small-$x$ TMD gluon distribution have been moderetely successful at describing the two-particle correlation data in $p+A$ colliisons [296, 297, 298, 299, 300], albeit with large theoretical uncertainties due to the naive LO approximation used. At LO, the CGC formalism predicts broadening of the away-side dihadron correlation. However, parton shower effects and the intrinsic $k_{T}$ of the fragmentation control the width of the correlation [301] and these effects are expected to reduce the broadening due to gluon saturation. However, recent studies [302] based on collinear factorization, nuclear PDFs and parton shower in the final state are also able to describe the majority of the dihadron and jet suppression effects in $p+A$ data at RHIC and the LHC. In such studies, the final state parton shower, accounting for the Sudakov resummation, is essential to reproduce the broadening of the away-side peak seen in the data.
@@
Second, double-parton scattering (DPS) can also at least partially obscure the interpretation of saturation effects, particularly in $p+A$ and even more substantially in $d+A$ collisions in which double-parton scattering is enhanced [304]. (Note that this DPS enhancement was not considered in Ref. [302].) One possibility would be to focus on $D^{0}\overline{D^{0}}$ correlations, as recently measured by LHCb [305]. The away-side peak also apparently disappears going from $p+p$ to $p+{\rm Pb}$ collisions at $p_{T}>2$ GeV. However, this transition can also be explained by nPDF + DPS [306], see Fig. 9. It would be interesting to investigate this observable from the CGC perspective. The LHCb measurement also differs from other correlation measurements, such as by STAR [307] by not normalizing the double yield ${dN}/{d^{2}p_{T}d^{2}q_{T}dy_{1}dy_{2}}$ by the single inclusive (trigger) yield. This is notable because shadowing or saturation can affect the single and double yields differently [302]. Direct photon-hadron/jet [308, 309, 310, 311, 312] or DIS lepton-hadron/jet [313, 314] correlations could prove valuable because they involve only one strongly interacting system in the final state.
@@
Figure 9: Dependence of $D^{0}\overline{D}^{0}$ production in $p+p$ and $p+\mathrm{Pb}$ collisions on the azimuthal separation $\phi$ between the $D^{0}$ and $\overline{D}^{0}$ measured by LHCb [305]. The calculations are based on collinear factorization [306]. The blue curve is a PYTHIA result with multiparton interactions turned off, thus only accounting for single-parton scattering. The red curve is a linear combination of the PYTHIA single particle prediction and a flat pedestal due to double-parton scattering. From Ref. [315].
@@
Finally, fragmentation of the partons to hadrons typically results in low $p_{T}$ hadrons produced from higher $p_{T}$ partons. The final state hadron momentum depends on the $z$, the ratio of the longitudinal momentum of the parton carried by the hadron. At midrapidity at LHC energies, $z\lesssim 0.5$ [316] while, away from midrapidty the average $z$ is higher because of the more restricted phase space for parton radiation at high rapidity. Thus, depending on the kinematics, the underlying partonic process may be in a kinematic region outside the saturation regime or even outside the TMD domain of validity. In this respect, $D^{0}\overline{D}^{0}$ correlations should be advantageous because the $\langle z\rangle\sim 0.7$ in $D$ meson production, see e.g. Fig. 4 of Ref. [317], due to different $z$ dependence of the charm fragmentation functions. Thus hadronization does not blur the underlying parton kinematics for $D$ mesons as much as it does for pions. A promising alternative, albeit challenging from the experimental point of view and not immune to DPS, is low $p_{T}$ back-to-back dijet production. See the following discussion as well as the $p+A$ studies in Refs. [318, 319, 320].

### v2

@@ III.1 Nuclear Structure Functions and nPDFs
In 1981, the EMC collaboration measured the nuclear modification of structure functions in deep inelastic scattering [1]. Due to the relation between structure functions and parton distribution functions at leading order and leading twist,
@@
these data revealed, for the first time, that the PDFs of a proton bound within a nucleus (nPDFs) differ from that of a free proton: $f_{i}^{p/A}(x,Q^{2})\neq f_{i}^{p}(x,Q^{2})$, where $f_{i}^{p/A}$ represents the PDF of a flavor $i$ parton inside a proton bound in a nucleus. The nuclear modification factor due to this effect is defined as $R_{i}^{A}\equiv f_{i}^{p/A}/f_{i}^{p}$. [^5] A slightly different definition is often adopted, $R_{i}^{A}\equiv\left(Zf_{i}^{p/A}+Nf_{i}^{n/A}\right)/\left(Zf_{i}^{p}+Nf_{i}^{n}\right)$, where $Z(N)$ is the number of protons (neutrons) in nucleus $A$, and $f_{i}^{n/A}(f_{i}^{n})$ are the bound (free) neutron PDFs. This definition takes into account the fact that the proton and neutron distributions in nuclei are always averaged and cannot be accessed separately. Since then, four main $x$ regions have been identified: at $x\gtrsim 0.7$, Fermi motion dominates with $R^{A}(x)\gtrsim 1$; in the range $0.3\lesssim x\lesssim 0.7$, the EMC region is characterized by $R^{A}(x)\lesssim 1$; between $0.1\lesssim x\lesssim 0.3$, antishadowing leads to $R^{A}(x)\gtrsim 1$; and finally, for $x\lesssim 10^{-1}$, shadowing is observed with $R^{A}(x)\lesssim 1$. Global analyses of nPDFs have been performed by the DSSZ [2], EPPS [3, 4], KSTSG [5], nCTEQ [6, 7], nNNPDF [8], and TUJU [9] collaborations. The latest analyses also include data from the LHC. For a recent review, see Ref. [210].
@@
  • Electroweak boson production: Production of $W$ and $Z$ bosons in $p+A$ collisions, $p+A\to Z,W+X$, measured at the LHC at $\sqrt{s_{NN}}=5.02$ and $8.16$ TeV probe quark nPDFs and, to some extent, the gluon nPDF in the range $10^{-3}\lesssim x\lesssim 10^{-1}$ in the shadowing and antishadowing regions at scale $Q^{2}\sim M_{W,Z}^{2}+p_{T}^{2}\gtrsim 6\times 10^{3}\,{\rm GeV}^{2}$ [219, 220, 221, 49, 46, 50, 52].
@@
The phenomenological extraction of the nuclear modification factor $R^{\text{nPDF}}$ critically depends on the data used in the global analyses. Earlier nPDF determinations relied primarily on DIS processes [227, 228, 229], where the initial lepton interacts purely via quantum electrodynamics (QED) with the nucleus, making them ideal for probing the quark nPDFs. More recently, data in $h+A$ collisions have been incorporated to extend the range of $x$ explored and access the gluon nPDFs more directly.
Figure 5 presents a comparison of recent quark and gluon nPDFs at $Q^{2}=10$ GeV2. The EPPS21 analysis [4] incorporates the $D$ production data in their global analysis, providing a strong constraint on $R_{g}^{\text{nPDF}}$ at $x\lesssim 10^{-2}$. Their result agrees with the nCTEQ15WZ analysis [230] but deviates from NNPDF2.0 [231], which does not include these data. These observations suggest that the nPDF evaluations remain highly sensitive to $h+A$ data used in the analyses.
@@
Including the LHC $p+{\rm Pb}$ measurements, the variety and the kinematical reach of the data used in nPDF analyses have grown significantly. It has therefore become increasingly possible to look for deviations from collinear factorization. For example, some of the nuclear effects seen in heavy flavor production in $p+{\rm Pb}$ collisions could be due to dynamical effects such as saturation [232], partonic energy loss [233], or collective effects. It is important to consider to what extent these effects are currently being incorporated in the nPDFs and how to potentially distinguish their ”non-factorizable” contribution. However, if the heavy flavor data are left out of the analysis, the small-$x$ gluon nPDF would be largely unconstrained and, as a result, the uncertainties on the baseline upon which these novel effects are supposed to be observed become prohibitively large. In other words, the baseline uncertainties overwhelm the effect one is trying to discover. In principle, this dilemma can be resolved if the remaining data are sufficiently precise and diverse, which is currently not the case. Because the factorizable effects are universal, the presence of effects beyond the nPDF formalism must eventually manifest as the inability to reproduce all the data with a single set of nPDFs in a statistically satisfactory way. It is thus useful to also view the global nPDF analyses as a search of physics beyond collinear factorization and not something that ignores or downplays them.
@@ III.2 Saturation effects in the small-$x$ regime
At small $x$, the gluon density inside hadrons grows rapidly, leading to nonlinear effects that are not captured by traditional linear QCD evolution equations such as DGLAP [167, 168] or BFKL [234, 235, 236, 237]. This phenomenon, known as gluon saturation, implies that the growth of the gluon distribution eventually slows down and reaches a limiting behavior. To describe this high-density regime, the CGC effective field theory has been developed. It models a high energy hadron as a coherent ensemble of soft gluons with large occupation numbers. By incorporating nonlinear gluon recombination effects, the CGC provides a more accurate and unified description of high-energy scattering processes. These nonlinear dynamics are characterized by the saturation scale $Q_{s}^{2}(x)\sim A^{1/3}\,x^{-\lambda}$, which grows as a power law with decreasing $x$. At sufficiently small $x$ (i.e., at high energies), the saturation regime becomes weakly coupled due to the semi-hard scale $Q_{s}\gg\Lambda_{\text{QCD}}$, making it amenable to perturbative QCD calculations within the CGC framework.
Figure 6 illustrates the current kinematic phase space for existing DIS and Drell-Yan data from nuclei heavier than iron, as a function of $x$ and $Q^{2}$. The saturation lines $Q^{2}(x)=Q_{s}^{2}(x)$ demarcate the transition between the linear and nonlinear regimes. The optimal region for studying these nonlinear effects is at $x\lesssim 10^{-2}$ and $Q^{2}\sim 1\text{ GeV}^{2}$. The kinematic reach of the future EIC, which will significantly extend access to this region, is also shown.
@@ III.3 Target universality from DIS to $p+A$
On the other hand, in the kinematic regime where the center-of-mass energy is very large and $Q^{2}$ is fixed but not too large, typically of the order of the saturation scale, collinear factorization is expected to break down. In this domain, the notion of weakly-interacting partons underlying parton distribution functions becomes inadequate. Instead, one must account for the nonlinear dynamics induced by large gluon densities and resum the large logarithms of the center-of-mass energy $\sqrt{s}$; specifically, $\ln(s/Q^{2})$ in DIS and $\ln(\sqrt{s}/p_{T})$ in hadronic collisions; which are commonly written as logarithms of $1/x$. The CGC formalism, where the coupling remains weak, but high-density effects are resummed to all orders, satisfies this requirement.
@@
where $x_{T},y_{T},z_{T}$ and $r_{T}$ are transverse coordinates. In multiparticle production in DIS and in proton-nucleus collisions, all other correlators are suppressed by powers of $N_{c}^{2}$ [239]. These correlation functions satisfy the JIMWLK evolution equations [240, 241] which encode the evolution of the correlators as function of $x$ once an initial condition at $x_{0}$ is specified. The large $Q^{2}$ and small $x$ regimes of QCD are unified in the double limit $Q^{2}\to\infty$ and $x\to 0$, where large virtuality and parton density effects interplay.
To leading order in the large $N_{c}$ approximation, the dipole and quadrupole correlators are the only objects that appear in physical cross sections within the CGC formalism. As such, they are universal quantities that can be probed by different processes, such as DIS and forward particle production in $p+A$ collisions [242, 243]. The universality of the cross sections is because the dipole and quadrupole correlators contain the small $x$ dynamics of the target independent of the projectile. It is then possible to extract the characteristics of the correlators from DIS and apply them to other processes, conferring the CGC formalism with predictive power. This is analogous to the collinear factorization framework, which is valid at high $Q^{2}$ but is expected to break down at very small $x$. Moreover, the universality of the Wilson line correlators mirrors that of the gluon distribution function governed by DGLAP evolution in the collinear factorization approach.
There have been sustained efforts to investigate higher-twist corrections to collinear factorization, particularly those enhanced by the size of the nucleus, scaling as $A^{1/3}$ [244, 150, 245, 246, 247]. [^6] Target mass corrections are important for understanding higher-twist effects at large $x$. These corrections can also be systematically included within the collinear factorization framework [248]. More recently, analogous studies have been initiated in the high-energy (small $x$) regime, focusing on subeikonal, energy suppressed contributions which are important for both unpolarized [249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260] and polarized observables [261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271], as well as for connecting the collinear and high-energy (small $x$) frameworks [272, 273, 274, 275, 276, 277, 278, 279].
These efforts motivate the search for observables that can be commonly accessed in both DIS and proton–nucleus collisions, enabling a consistent characterization of the small $x$ structure of nuclei for a variety of probes.
@@ III.4 Observables
In this section, we explore several key observables that may distinguish between nuclear parton distribution function and gluon saturation effects. So far, the extraction of nPDFs at small $x$ has not taken other physical effects into account. Determining the relative contribution of each effect is crucial, as there are limited small $x$ data available, e.g. at $x\sim 10^{-5}$ in $h+A$ collisions, as discussed in Sec. III.1.
III.4.1 Structure functions in DIS
CGC-based observables have a characteristic $x$ dependence. While the theoretical uncertainties associated with the perturbative expansion are, in principle, controllable, additional uncertainties arise from modeling the nonperturbative inputs required to initialize the small $x$ evolution. The initial conditions are typically parametrized and fitted to experimental data at an intermediate $x$ value, $x_{0}\sim 10^{-2}$ [280, 281, 282, 283]. Once this initial condition is fixed, the CGC framework provides first principles predictions of the $x$ dependence of the structure functions $F_{2}$ and $F_{L}$, generally measured in DIS.
The EIC will cover a large kinematic range in $x$, about three orders of magnitude in the region where $x$ can be considered small, which should make it possible to distinguish between shadowing generated by the perturbative small $x$ evolution of the target via the CGC and DGLAP evolution of the $x$-dependent parameterization of shadowing models. The $x$ dependence of these two approaches is quite different. For example, DGLAP evolution does not generate any additional shadowing while, in the CGC, it is dynamically generated. Unfortunately, the $F_{2}$ structure function necessarily receives sizable contributions from nonperturbatively large dipoles [284] at small $x$, even at intermediate $Q^{2}$, making it difficult to distinguish perturbative contributions from nuclear shadowing. The longitudinal structure function $F_{L}$ is under much better theoretical control. However, determinations of $F_{L}$ are much more difficult because it requires measurements at different energies.
@@ III.4.2 Single and two-particle production in $e+A$ and $p+A$
Alongside the fully inclusive processes discussed in the previous section, saturation and shadowing also leave their imprints on more differential observables, such as single and double-inclusive particle production in DIS and forward production in proton-nucleus collisions.
@@
A promising production channel for observing saturation is the disappearance of the away-side peak in two-particle azimuthal correlations in $e+A$ or $p+A$ collisions at forward rapidity [286, 287, 288]. In the strict back-to-back limit, CGC calculations, expanded to leading power in the ratio between the transverse momentum imbalance and the relative dijet momentum, smoothly match to TMD factorization [289]. In particular, back-to-back dihadron production at LO probes the Weizsäcker-Williams gluon distribution while direct photon-jet correlations in forward $p+A$ collisions are sensitive to the dipole gluon TMD [289]. Beyond LO in the CGC, two-particle correlations also probe the sea quark TMD distribution [290]. This correspondence between CGC predictions and TMD factorization provides a powerful framework for addressing the question posed at the beginning of Sec. III, whether the suppression of the away-side peak is purely nonperturbative in origin (as encoded in the nonpertubative part at small transverse momenta of the TMDs or in the nPDF initial condition of Collins-Soper-Sterman evolution [291, 292, 293, 294]) or whether gluon saturation itself demonstrates the perturbative mechanism at work.
@@
There is, however, a complication to the previous general picture: when the two particles are back-to-back in the transverse plane, higher order QCD corrections, enhanced by large logarithms in the ratio of the relative transverse momentum to the transverse momentum imbalance, further decorrelate the away side peak by a vacuum-like Sudakov suppression factor (or Collins-Soper-Sterman evolution [291, 292, 293, 294] in the TMD language).
@@
While this complication makes interpretation of the cause of suppression more ambiguous, it is systematically included both in CGC calculations and in standard TMD factorization with increasing accuracy, enabling the search for saturation effects “on top of” the modifications caused by soft gluon radiation, present even in the absence of multiple soft scattering or nonlinear effects. The interplay between Sudakov resummation and gluon saturation or small $x$ evolution has been considered in Refs. [295, 204, 296, 297].
@@
Thus, in discussions of two-particle azimuthal correlations, it is useful to distinguish between broadening and suppression: saturation is expected to lead to both broadening (due to multiple scattering) and suppression (due to small $x$ radiation and nonlinear evolution) whereas radiative corrections associated with large-angle soft gluon emissions (the Sudakov effect) primarily result in broadening alone. The CGC approach provides a physical mechanism, gluon saturation, to generate the dependence of the nuclear suppression on the $p_{T}$ of the associated hadron. In the nPDF paradigm, in turn, the $p_{T}$ dependence enters through the average $x$ value probed (the smaller the $p_{T}$, the smaller the $x$), as well as through the value of the factorization scale, typically proportional the the hadron $p_{T}$. In addition, DGLAP evolution tends to reduce the nuclear effects at larger factorization scales.
@@
Numerical predictions based on the TMD framework supplemented by CGC models of the small $x$ TMD gluon distribution have been moderately successful at describing the two-particle correlation data in $p+A$ collisions [298, 299, 300, 301, 302], albeit with large theoretical uncertainties due to the naive LO approximation used. At LO, the CGC formalism predicts broadening of the away-side dihadron correlation. However, parton shower effects and the intrinsic $k_{T}$ of the fragmentation control the width of the correlation [303] and these effects are expected to reduce the broadening due to gluon saturation. However, recent studies [304] based on collinear factorization, nPDFs and parton showers in the final state are also able to describe the majority of the dihadron and jet suppression effects in $p+A$ data at RHIC and the LHC. In such studies, the final state parton shower, accounting for the Sudakov resummation, is essential to reproduce the broadening of the away-side peak seen in the data.
@@
Second, double-parton scattering (DPS) can also at least partially obscure the interpretation of saturation effects, particularly in $p+A$ and even more substantially in ${\rm d}+A$ collisions in which double-parton scattering is enhanced [306]. (Note that this DPS enhancement was not considered in Ref. [304].) One possibility would be to focus on $D^{0}\overline{D^{0}}$ correlations, as recently measured by LHCb [307]. The away-side peak also seems to disappear going from $p+p$ to $p+{\rm Pb}$ collisions at $p_{T}>2$ GeV. However, this transition can also be explained by nPDF + DPS [308], see Fig. 9. It would be interesting to investigate this observable from the CGC perspective. The LHCb measurement also differs from other correlation measurements, such as by STAR [309] by not normalizing the double yield ${dN}/{d^{2}p_{T}d^{2}q_{T}dy_{1}dy_{2}}$ by the single inclusive (trigger) yield. This is notable because shadowing or saturation can affect the single and double yields differently [304]. Direct photon-hadron/jet [310, 311, 312, 313, 314] or DIS lepton-hadron/jet [315, 316] correlations could prove valuable because they involve only one strongly interacting system in the final state.
@@
Figure 9: Dependence of $D^{0}\overline{D}^{0}$ production in $p+p$ and $p+\mathrm{Pb}$ collisions on the azimuthal separation $\phi$ between the $D^{0}$ and $\overline{D}^{0}$ measured by LHCb [307]. The calculations are based on collinear factorization [308]. The blue curve is a PYTHIA result with multiparton interactions turned off, thus only accounting for single-parton scattering. The red curve is a linear combination of PYTHIA single particle prediction and a flat pedestal due to double-parton scattering. From Ref. [317].
@@
Finally, fragmentation of the partons to hadrons typically results in low $p_{T}$ hadrons produced from higher $p_{T}$ partons. The final state hadron momentum depends on $z$, the ratio of the longitudinal momentum of the parton carried by the hadron to the initial parton momentum. At midrapidity at LHC energies, $z\lesssim 0.5$ [318], while, away from midrapidty the average $z$ is higher because of the more restricted phase space for parton radiation at higher rapidities. Thus, depending on the kinematics, the underlying partonic process may be in a kinematic region outside the saturation regime or even outside the TMD domain of validity. In this respect, $D^{0}\overline{D}^{0}$ correlations should be advantageous because $\langle z\rangle\sim 0.7$ in $D$ meson production (see e.g. Fig. 4 of Ref. [319]) because charm fragmentation function has a different $z$ dependence than those of light quarks. Thus hadronization does not blur the underlying parton kinematics for $D$ mesons as much as it does for pions. A promising alternative, albeit challenging from the experimental point of view and not immune to DPS, is low $p_{T}$ back-to-back dijet production. See the following discussion as well as the $p+A$ studies in Refs. [320, 321, 322].